[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Mon Oct 31, 2016

TAFB Shift Log nhc.tsbadmin at noaa.gov
Mon Oct 31 11:59:40 UTC 2016


TAFB Shift Log for Mon Oct 31 2016 07:59:37 GMT-0400 (EDT)

Shift: Day Shift

Pacific Forecaster: A Levine

Atlantic Forecaster: J Lewitsky

Surface Analyst: N Ramos

Technical Issues Summary:

GFE/NCP issues:

N-AWIPS issues:

Other issues: A/C issues last night (see previous log entry), still  
relatively warm in ops area. Light fixture still needs to be fixed over  
TAFB backup desk area.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161031/3fa14b78/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list